# Graph neural network surrogate for strategic transport planning

### v1

@@ 4.1 Comparison of GCNII, GATv2, GATv3 and GCN+GATv3
The augmented dataset is used. All the models in this experiment are evaluated using the average F1 score. For the GCNII, a grid search is performed over the number of GCNII layers $\in$ (5, 10, 50, 70), the size of each layer $\in$ (64, 256, 512), $\alpha$in (0.1, 0.4) and $\theta$in (0.5, 1.5). The best performing model on the validation dataset is based on 70 layers, each layer containing 512 hidden units, $\alpha$ = 0.1 and $\theta$ = 1.5. If not stated otherwise, the same parameters are used for all future GCNII experiments. This experiment further investigates whether the more expressive GAT models improve the predictions. In the preliminary analysis, it was observed that the GATv2 model had convergence problems when attempting to stack more than 5 layers, which is also supported by the observations in the literature [Chen et al., 2020]. Due to the average graph diameter of the dataset being 36, it is mathematically impossible for a shallow GATv2 model to correctly capture the underlying travel patterns. Thus, a GATv3 grid search is performed, with either 20 hidden layers and 512 hidden units or 40 hidden layers with 256 hidden units. Based on the validation dataset, the best configuration is 20 hidden layers with 512 hidden units. Finally, to explore whether the over-squashing and over-smoothing problems can be overcome, the combined GCNII and GATv3 model is investigated. Here, two configurations are explored, either 5 or 10 layers of GATv3 (using the zone connected graph), whilst the following 60 layers are GCNII layers using the full network graph. The best model is found to contain 10 GATv3 layers.
4.2 Hierarchical regression
Classification has less data parameters to tune, thus making it often more stable than regression, but does not produce exact values. To combine both the approaches, the hierarchical regression setup is investigated, which to the best of the authors’ knowledge, does not exist in literature. The core idea is to first predict the bucket in which the transport usage should be. In the next step, given the bucket prediction, a value is outputted between -1 and 1, so as to be able to predict exact values for any bucket by applying normalization. Using the bucket and the regression prediction, it is then possible to determine the exact value being predicted. The core idea of the approach is to combine the stability of classification with the precision of regression.
The input data setup is identical to the previous experiment, i.e., using the augmented dataset and direct graph conversion. The output data is transformed to have two targets: the bucket class and the scaled value within that bucket. The configuration of the output data is thus dependent on the buckets selected. A number of bucket configurations are explored, using the same manual search strategy as employed in the previous regression experiments. The tested bucket groups are the following:
  • Bucket T1 - Buckets [0, 10), [10, 500), $>$500 veh/h - investigated due to their success in the previous experiments
  • Bucket T2 - Buckets [0, 100), [100, 500), $>$500 veh/h - inspired by the first bucket configuration, but to have more equal sized buckets, without skewing the sample distribution too much
  • Bucket T3 - Buckets [0, 10), [10, 100), [100, 200), [200, 500), [500, 1000), [1000, 2000), $>$2000 veh/h - to exploit the high number of data points available in the first few buckets
  • Bucket T4 - Buckets going in steps of 100 veh/h from 0 to 4500 - covers all the data points with equidistant buckets
As base GNN model, the GCNII is taken with 70 GCNII layers, each with 512 hidden units, $\alpha=0.1$ and $\theta=1.5$. All the models are trained for 300 epochs, taking around 8-12 hours per training run. To implement the hierarchical regression, the classification is performed by two feed forward layers followed by a softmax activation function. The argmax is taken for this output, and then an embedding vector of size 32 is retrieved for this class, inspired by a similar mechanism in natural language processing [Bahar et al., 2020]. The embedding vector is then concatenated to the output of the GCNII layers, and is used by two feed forward layers to produce the output of the link using the tanh activation function, for determining the exact value. A formal definition is as follows:
[…]
$g^{regr}(x)=$ $\,\sigma_{tanh}\left(N_{ff}\left(\sigma_{relu}\left(N_{ff}\left(x\right)\right)\right)\right)$ (4)
$g^{class}(x)=$ $\,\operatorname*{arg\,max}_{k}\left[\sigma_{softmax}\left(N_{ff}\left(\sigma_{relu}\left(N_{ff}\left(x\right)\right)\right)\right)_{k}\right]$ (5)
where,
$g(x)_{i}$ is the final output of the model for link i
$g^{regr}(x)_{i}$ is the output of the regression model for link i
$g^{class}(x)_{i}$ is the output of the classification model for link i $N_{ff}$ is a single feed forward layer
[…]
[;][;][ ; ] is the concatenating operator
[…]
$N_{embedding}$ function is a differentiable operator that takes a transport usage bucket ID as input and retrieves a differentiable vector out of memory
The aim of the $N_{embedding}$ is to make the neural network automatically learn to represent each bucket with its own custom vector, without having to explicitly define this vector. An illustration is provided in Figure 2. Training for the classification is done using cross entropy, whilst the regression component is trained using MSE. The losses are added together to form the final loss. Different weightings of the loss functions did not show any difference in performance. To stabilize the training, true labels are used as input to the embedding step, but only during the training of the model. GraphNorm was attempted to be used, but resulted in no convergence for the models. Additionally, residual final layers could not be employed as the use of them lead to GPU memory issues.
Figure 2: Illustration of the hierarchical regression model.
4.3 Fine grained classification
@@
To further exploit the stability of classification, this experiment is conducted to explore fine grained classification. Furthermore, a method is proposed to convert the classification results into real values. Comparing the F1 score does not make sense when using different buckets, as the the comparison becomes skewed toward the model configurations with fewer buckets. Thus, it makes more sense to use the following metrics: $R^{2}_{\,\geq 10}$ and $MAE_{\,\geq 10}$. However, in classification, only the probabilities are provided as outputs of the model. To convert the probabilities of buckets into a real value, a small transformation procedure is applied using the expectation operator from statistics. Initially, a uniform distribution across each bucket is assumed and the expectation is taken, resulting in the mean value of the bucket. Then, the expectation over all buckets is applied, using the output of the model as the discrete probability distribution. The result is that the model outputs real values in veh/h. Formally,
@@
4.3.1 Optimal Bucket Strategy
@@
4.3.2 Use of additional synthetic data
@@
The final experiments examine how the GATv3 and GCN+GAT models perform using the training data supplemented by the synthetic samples. The data setup is identical to the previous experiment, but only using the Buckets-nl54 configuration. The GCNII is used with 70 layers, 512 hidden units, $\alpha=0.1$, $\theta=1.5$, 3 residual feed forward layers, GraphNorm after each convolution operation and a dropout of 25%. The GATv3 model contains 20 layers each with 512 hidden units, based on 2 attention heads performing the averaging merging strategy, 3 residual feed forward layer, GraphNorm and dropout with 25%. Finally, the GCN+GAT model is configured with 10 zonal GATv3 layers, each with 128 hidden units, 2 attention heads being averaged together. Following that, 60 GCNII layers are used, each with 512 hidden units, $\alpha=0.1$, $\theta=1.5$ and 3 residual feed forward layers. GraphNorm and dropout of 25% is used for each convolution layer. Note, each training run takes around 40 hours.
@@ 5 Experimental results
5.1 Evaluation of different GNN architectures (GCNII, GATv3, GCN+GATv3) and hierarchical regression
Different GNN models are compared to find the optimal model setup. The F1 score of all the models can be found in Table 1. It is observed that the GCNII model performs significantly better than all baselines. Note that the accuracy of the model is at 85%, and most errors are observed in classifying the [10 veh/h, 500 veh/h) bin. For the bin of [0 veh/h, 10 veh/h), an accuracy of 95% is achieved. On a different note, the configuration containing both the largest number of layers and largest layer size in the GCNII performs best. The best performing model is the GATv3, followed closely by the GCN+GAT model, and then the GCNII model. All three GNNs perform similarly to each other, and significantly better than the baselines.
@@ 5.1 Evaluation of different GNN architectures (GCNII, GATv3, GCN+GATv3) and hierarchical regression
With regards to the model performance, the three GNN models perform well and have only slight variations. First, this implies that deep GNNs are a suitable candidate to further explore for a surrogate model for the four step model. Secondly, it implies that the exact model does not make a major difference. For future experiments, a GCNII can be used for initial experiments, and other GNN models can be tested, once a potential stable problem setup is found. Looking at the GCN+GAT model, it has similar performance to the other models. A hypothesis on this is that either it does not solve the over-squashing problem, or the issue does not exist for this initial task. Further research into this should be examined in future experiments.
Looking at the hierarchical regression, the best model using “Bucket 3” option reaches only 175.8 $MAE_{\,\geq 10}$, significantly worse than the best regression model from Narayanan et al. [2024]. Even though the hierarchical regression models perform better than the baselines, they do not come close to the best model from the classical regression experiments. The subpar performance found in hierarchical regression may be due to the fact that the task is too complex as a problem for the model experimented here. Deeper models with more parameters could potentially be a solution, but there needs to be further research on this topic.
@@ 5.2 Evaluation of fine grained classification
In this section, the results of the classification performed on fine grained buckets are presented. By applying the double expectation operator, it is possible to convert probabilities to real values, with all details about the setup found in Section 4.3. With regards to the optimal bucket strategy, the resulting evaluation metrics of the relevant experiments are shown in Table 2. It can be seen that all reach good results, when compared to the baselines. However, only GCNII - Buckets-e90 slightly surpasses the $MAE_{\,\geq 10}$ score of the best regression model. Between the best models of the linear and non-linear bucketing, the latter does not surpass the former.
The two best performing bucket configurations, nl54 and e90, are then trained from scratch using additional synthetic training data. Both the models with the extra data perform significantly better than the baselines, as can be seen in Table 2, and have an 16% better performance than the models without the extra data. However, the best non-linear bucketing model has a slightly better performance than the best linear bucketing model, which is in contrast to the ones without the extra data. Looking more closely at the differences between the models “GCNII - Buckets-e90 - extra data” and “GCNII - Buckets-nl54 - extra data”, it is found that the former predicts high values poorly. As seen in Figure 3, all true values above around 2500 veh/h get assigned values around 2600 veh/h. On the other hand, the model with non-linear 54 buckets continues to predict proportionally high values, as shown in Figure 4. The effect comes probably from the high number of classes having few data points for training in the case of equidistant bucketing.
Table 2: Results of fine grained classification on validation dataset
Model Validation $R^{2}_{\,\geq 10}$ Validation $MAE_{\,\geq 10}$
Mean Regressor 0.00 394.8
Random Forest 0.00 373.1
MLP 0.00 371.6
GCNII - Regression Best 0.84 131.1
GCNII - Buckets-e23 0.79 145.4
GCNII - Buckets-e45 0.78 140.1
GCNII - Buckets-e90 0.81 129.6
GCNII - Buckets-e180 0.78 139.5
GCNII - Buckets-e450 0.78 137.8
GCNII - Buckets-nl38 0.79 136.1
GCNII - Buckets-nl54 0.79 133.0
GCNII - Buckets-e90 - extra data 0.85 114.8
GCNII - Buckets-nl54 - extra data 0.86 111.4
@@
Compared to the regression setting, the fine grained classification returned more consistent results, probably as fewer steps are required. Such behaviour is especially desirable as it requires less configuration searching and reduces the overall model development time. Furthermore, extra synthetic data has shown to make a significant improvement in the performance of the models. To validate the result with more GNNs, GATv3 and GCN+GAT models are trained. The evaluation metrics of the experiment runs can be found in Table 3. The best performing model is the GCNII with extra data, closely followed by the GCN+GAT model. But with GATv3, the performance is comparatively poor. Narayanan et al. [2024] observed the existence of positive correlation between the graph size and the errors in the regression problem, meaning that the GNNs are not able to generalize well to larger networks. The same pattern continue to exist even in the fine grained classification problem, as depicted in Figure 5. On a different note, the improvements observed by increasing the training data is consistent with different GNN models. This implies that more training data is required for better predictions, as is the usual case for the deeper models.
@@
GCN+GAT - Buckets-nl54 - extra data 0.84 120.0
Figure 5: $MAE_{\,\geq 10}$ with respect to graph size, i.e. number of nodes for the trained GCN+GAT model on the test dataset. Each point is the average absolute error of a single graph, with the errors coming from incorrect link transport usage prediction. Highly similar plots are observed for all GNNs.
@@ 6 Conclusions
In this study, we embarked on an exploration of surrogate modelling for strategic transport planning, leveraging advanced Graph Neural Network (GNN) architectures. Specifically, we compared the established Graph Convolution Network (GCN) with the more recent and expressive Graph Attention Network (GAT). Additionally, we introduce a novel variant of GAT (namely GATv3) tailored to address the issue of over-smoothing present in graph-based models. Furthermore, we investigate (i) the synergistic potential of combining both GCN and GAT architectures, (ii) the realm of hierarchical regression setups integrating classification and regression tasks and (iii) the realm of fine-grained classification combined with a method to convert its output to precise values, to avoid the issues associated with the complexity of regression and hierarchical setups. Finally, a synthetic data generator is implemented to study the impact of increasing the training data.
Our findings shed light on the performance of various model architectures across distinct tasks, providing valuable insights for future research and application in this domain. Our results underscore the effectiveness of the newly proposed Graph Attention Network (GATv3) in the classification task. Thus, it is possible to develop deep GATs using the GATv3 formulation, which is relevant for problems wherein the feature sharing between distant nodes is important. Furthermore, the superior performance of the new GAT, when compared to existing GATs and the GCN models, highlights its potential as a formidable tool for capturing more intricate patterns within transportation networks. This signifies a significant step forward in refining surrogate models for strategic transport planning.
The combined GAT and GCN model, while exhibiting a marginal improvement over the standalone GCN model, was unable to surpass the performance of the new GAT. This observation suggests that while hybrid architectures hold promise, there is a need for further investigation to increase their synergistic potential. Our exploration into hierarchical regression, which aimed to combine categorical and continuous aspects of transportation dynamics, yielded suboptimal results. This indicates that deeper analysis are warranted to understand the challenges and potential solutions in this specific modelling context. Surprisingly, in the fine-grained classification task with the provision of additional training data (produced through the propsed synthetic data generator), the Graph Convolution Network (GCN) emerged as the top performer. This contrasts with the performance of the new GAT and the combined GAT+GCN models, which, while competitive, fell slightly short. This outcome prompts consideration of potential factors, such as the problem type and the impact of expanded dataset.
To sum up, the answer to whether fine-grained classification can perform well is yes. Additionally, it is observed to be more stable with respect to the model hyperparameters than regression. With an extended training dataset, further performance improvements are reached. Therefore, exploring strategies to augment the dataset could lead to further advancements. Nevertheless, a detailed analysis shows that the positive correlation between network size and error continues to be an issue, which requires further exploration in future. In conclusion, this research represents a significant stride towards refining GNN based surrogate models for strategic transport planning. The insights gleaned from this study pave the way for several promising avenues of future research. The demonstrated superiority of the new GAT in the classification task and the unexpected dominance of GCN in fine-grained classification underscore the need for continued exploration and refinement of GNN architectures in this context. Furthermore, the novel GAT architecture and the modelling setups (such as the hierarchical regression and the fine-grained classification with the possibility of obtaining the exact values) proposed in this study could be explored for other transportation problems. We anticipate that this work will serve as a valuable foundation for future endeavors in enhancing the accuracy and applicability of GNN-based surrogate models in complex urban planning scenarios.

### v2

@@ 3.3 Experimental setup
The augmented dataset is used. All the models in this experiment are evaluated using the average F1 score. For the GCNII, a grid search is performed over the number of GCNII layers $\in$ (5, 10, 50, 70), the size of each layer $\in$ (64, 256, 512), $\alpha$in (0.1, 0.4) and $\theta$in (0.5, 1.5). The best performing model on the validation dataset is based on 70 layers, each layer containing 512 hidden units, $\alpha$ = 0.1 and $\theta$ = 1.5. If not stated otherwise, the same parameters are used for all future GCNII experiments. This experiment further investigates whether the more expressive GAT models improve the predictions. In the preliminary analysis, it was observed that the GATv2 model had convergence problems when attempting to stack more than 5 layers, which is also supported by the observations in the literature [Chen et al., 2020]. Due to the average graph diameter of the dataset being 36, it is mathematically impossible for a shallow GATv2 model to correctly capture the underlying travel patterns. Thus, a GATv3 grid search is performed, with either 20 hidden layers and 512 hidden units or 40 hidden layers with 256 hidden units. Based on the validation dataset, the best configuration is 20 hidden layers with 512 hidden units.
3.3.2 Fine grained classification
3.3.2.1 Overview of the setup
@@
To further exploit the stability of classification, this experiment is conducted to explore fine grained classification. Furthermore, a method is proposed to convert the classification results into real values. Comparing the F1 score does not make sense when using different buckets, as the the comparison becomes skewed toward the model configurations with fewer buckets. Thus, it makes more sense to use the following metrics: $R^{2}_{\,\geq 10}$ and $MAE_{\,\geq 10}$. However, in classification, only the probabilities are provided as outputs of the model. To convert the probabilities of buckets into a real value, a small transformation procedure is proposed using the expectation operator from statistics. Initially, a uniform distribution across each bucket is assumed and the expectation is taken, resulting in the mean value of the bucket. Then, the expectation over all buckets is applied, using the output of the model as the discrete probability distribution. The result is that the model outputs real values in veh/h. Formally,
@@
3.3.2.2 Optimal Bucket Strategy
@@
3.3.2.3 Use of additional synthetic data
@@
The final experiments examine how GATv3 model perform using the training data supplemented by the synthetic samples. The data setup is identical to the previous experiment, but only using the Buckets-nl54 configuration. The GCNII is used with 70 layers, 512 hidden units, $\alpha=0.1$, $\theta=1.5$, 3 residual feed forward layers, GraphNorm after each convolution operation and a dropout of 25%. The GATv3 model contains 20 layers each with 512 hidden units, based on 2 attention heads performing the averaging merging strategy, 3 residual feed forward layer, GraphNorm and dropout with 25%.
@@ 4 Experimental results
4.1 Evaluation of different GNN architectures (GCNII and GATv3)
Different GNN models are compared to find the optimal model setup and a classification task is used for this. The choice of classification as the primary task for comparing different GNN architectures was driven by methodological and practical considerations. Classification provides a simpler and more controlled experimental setting, allowing for a direct comparison of architectures without the added complexities of continuous target prediction. Moreover, classification tasks are generally more stable in terms of optimization and convergence, reducing the risk of confounding factors such as sensitivity to hyperparameters, which are more prominent in regression. Starting with classification also allowed for a structured experimental progression - by first assessing how well each architecture performs on a simpler task, we could identify potential strengths and weaknesses before moving to the more complex problem. Additionally, classification outputs are inherently more interpretable, enabling easier diagnostic analysis to understand model behavior before tackling a more nuanced regression setting. Overall, choosing classification as the initial task ensured a robust and systematic comparison of GNN architectures, minimizing external sources of variability and establishing a solid basis for the subsequent experiments.
In a preliminary analysis, it was observed that the GATv2 model had convergence problems when attempting to stack more than 5 layers, which is also supported by the observations in the literature [Chen et al., 2020]. Due to the average graph diameter of the dataset being 36, it is mathematically impossible for a shallow GATv2 model to correctly capture the underlying travel patterns. Hence, the results are shown only for GCNII and GATv3 GNN models. The F1 score of the models can be found in Table 1. It is observed that the GCNII model performs significantly better than all baselines. Note that the accuracy of the model is at 85%, and most errors are observed in classifying the [10 veh/h, 500 veh/h) bin. For the bin of [0 veh/h, 10 veh/h), an accuracy of 95% is achieved. On a different note, the configuration containing both the largest number of layers and largest layer size in the GCNII performs best. Nevertheless, the best performing model is the GATv3.
@@ 4.2 Evaluation of fine grained classification
In this section, the results of the classification performed on fine grained buckets are presented. By applying the double expectation operator, it is possible to convert probabilities to real values, with all details about the setup found in Section 3.3.2. Different bucket strategies reach good results, when compared to the baselines. However, only GCNII - Buckets-e90 slightly surpasses the $MAE_{\,\geq 10}$ score of the best regression model. Between the best models of the linear and non-linear bucketing, the latter does not surpass the former.
The two best performing bucket configurations, nl54 and e90, are then trained from scratch using additional synthetic training data. Both the models with the extra data perform significantly better than the baselines and have an 16% better performance than the models without the extra data. However, the best non-linear bucketing model has a slightly better performance than the best linear bucketing model, which is in contrast to the ones without the extra data. Looking more closely at the differences between the models “GCNII - Buckets-e90 - extra data” and “GCNII - Buckets-nl54 - extra data”, it is found that the former predicts high values poorly. As seen in Figure 2, all true values above around 2500 veh/h get assigned values around 2600 veh/h. On the other hand, the model with non-linear 54 buckets continues to predict proportionally high values, as shown in Figure 3. The effect comes probably from the high number of classes having few data points for training in the case of equidistant bucketing.
@@
Compared to the regression setting, the fine grained classification returned more consistent results, probably as fewer steps are required. Such behaviour is especially desirable as it requires less configuration searching and reduces the overall model development time. Furthermore, extra synthetic data has shown to make a significant improvement in the performance of the models. The evaluation metrics of the experiment runs can be found in Table 2. The best performing model is the GCNII with extra data, and the performance is comparatively poor with GATv3.
Narayanan et al. [2024] observed the existence of positive correlation between the graph size and the errors in the regression problem, meaning that the GNNs are not able to generalize well to larger networks. The same pattern continue to exist even in the fine grained classification problem, as depicted in Figure 4. On a different note, the improvements observed by increasing the training data is consistent with different GNN models. This implies that more training data is required for better predictions, as is the usual case for the deeper models.
@@
Figure 4: $MAE_{\,\geq 10}$ with respect to graph size, i.e. number of nodes for the trained GNN model on the test dataset. Each point is the average absolute error of a single graph, with the errors coming from incorrect link transport usage prediction. Highly similar plots are observed for all GNNs.
@@ 5 Conclusions
In this study, we developed advanced Graph Neural Network (GNN) architectures as surrogate models for strategic, long-term transport planning. Our work involved a comparative analysis between Graph Convolutional Networks (GCNs) and the more expressive Graph Attention Networks (GATs). To address the over-smoothing issue that limits the depth of GAT models, we introduced GATv3, a novel variant incorporating residual connections. Additionally, we proposed a fine-grained classification approach (along with a method to convert its output into precise values) that balances predictive stability with numerical precision, mitigating the challenges posed by complex regression tasks. To further improve model performance, we developed a synthetic data generator to expand training datasets without overfitting.
Our findings provide important insights into the performance of different model architectures, offering valuable guidance for future research and practical applications in the field. Notably, GATv3 demonstrated strong classification capabilities. The GATv3 formulation enables the development of deeper GAT models, making it particularly well-suited for classification-based transportation applications, where feature sharing across distant nodes is crucial. The results from the study also confirm that fine-grained classification is more stable with respect to model hyperparameters compared to regression tasks. Interestingly, in this task, GCNII emerged as the top performer when supplemented with additional synthetic training data, surpassing GATv3. All these suggest that problem type, task complexity, and dataset expansion play crucial roles in model performance. Despite these advancements, a detailed analysis reveals a linear relationship between network size and residual error, likely due to error propagation in message passing. As networks grow, errors accumulate and spread across nodes, necessitating further research into mitigating this effect in larger GNNs.
Overall, this study marks a significant step forward in refining GNN-based surrogate models for transport planning. The proposed GATv3 architecture and fine-grained classification framework offer promising new directions for broader transportation applications. On the one hand, GATv3 appears well-suited for classification-based transportation applications, such as section control and congestion warning systems, which require a higher degree of differentiation among neighboring links and feature sharing across distant nodes. On the other hand, GCNII is suited for the development of surrogates for travel demand modelling. Moving forward, future research should focus on reducing error propagation in deep GNNs, optimizing data augmentation strategies, and further refining GNN architectures for complex urban mobility scenarios.